40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 4753     Entry time: Fri May 20 03:01:17 2011
Author: kiwamu 
Type: Update 
Category: LSC 
Subject: locking status 

(PRMI locking)

Since REFL11 has gone I tried locking the PRMI with combination of REFL55 and AS55.

Without any pain the lock of PRMI was achieved successfully. AS55 was used to sense MICH and REFL55 was used for PRC.

 

(scripting)

Additionally I was modifying several scripts which are invoked from C1IFO_CONFIGURE.adl. Some details about the scripts will be uploaded on the wiki later.

An important thing is that now we are able to use the "restore" commands for the Y arm, X arm, Michelson and PRM locking.

The scripts will automatically acquire the lock of each DOF.  The image below is just a screen shot of the medm screen where you can call the scripts.

Screen_shot_2011-05-20_at_2.50.00.png

 

( Still to do)

   * PRM actuator response measurement

   * PRC noise budget

   * MICH-PRC actuator decoupling

ELOG V3.1.3-